OpenStudy (anonymous):

express in partial fractions: (2x^4 -3x^3+4x-2)/(x^2-x+1)²(x-1)

OpenStudy (anonymous):

\[\frac{ (2x^4 -3x^3+4x-2) }{ (x^2-x+1)² (x-1)}\]

OpenStudy (anonymous):

or

OpenStudy (anonymous):

\[\frac{ (2x^4 -3x^3+4x-2) }{ (x^2-x+1)² }(x-1)\]

OpenStudy (anonymous):

what's your Question ?

OpenStudy (anonymous):

\[\frac{2x^4-3x^3+4x-2}{(x^2-x+1)^2(x-1)}=\frac{Ax+B}{x^2-x+1}+\frac{Cx+D}{(x^2-x+1)^2}+\frac{E}{x-1}\]